the modulus or absolute value of a complex numbers:

if z2=4+3i and z2=-2+3i what is /z1/ and /z2/?


Answer by stanbon(29500)

if z2=4+3i and z2=-2+3i what is /z1/ and /z2/?
--------------------------------------
===============================================
|4+3i| = sqrt(4^2 + 3^2) = 5
|-2+3i| = sqrt(2^2 + 3^2) = sqrt(13)
====================================
Cheers,
Stan H.